The market price is ​$800 for a 19​-year bond ​($1,000 par​ value) that pays 11 percent annual​ interest, but makes interest payments on a semiannual basis ​(5.5 percent​ semiannually).

What is the​ bond's yield to​ maturity?
Answer #1

Number of periods = 19 * 2 = 38

Coupon = (0.055 * 1000 = 55

Yield to maturity = 14.04%

Keys to use in a financial calculator: 2nd I/Y 2, FV 1000, N 38, PMT 55, PV -800, CPT I/Y